Let logbA = 2.487 and logbB = 0.392. Find logb(A/B).
Answer by kapilsinghi(68)

Let logbA = 2.487 and logbB = 0.392. Find logb(A/B).
logb(A/B) = logb(A) - logb(B)
= 2.487 - 0.392
= 2.095

Regards
Kapil
You can send your questions at kapilsinghi123@gmail.com